# Asymptotics of the Hankel determinant and orthogonal polynomials arising from the information theory of MIMO systems

### v1

@@ Abstract
We consider the Hankel determinant and orthogonal polynomials with respect to the deformed Laguerre weight
$w(x;t)={x^{\alpha}}{\mathrm{e}^{-x}}{(x+t)^{\lambda}},\qquad x\in\mathbb{R}^{+},$
with parameters $\alpha>-1,\;t>0$ and $\lambda\in\mathbb{R}$. This problem originates from the information theory of single-user multiple-input multiple-output (MIMO) systems studied by Chen and McKay [IEEE Trans. Inf. Theory 58 (2012) 4594–4634]. By using the ladder operators for orthogonal polynomials with general Laguerre-type weights, we obtain a system of difference equations and a system of differential-difference equations for the recurrence coefficients $\alpha_{n}(t)$ and $\beta_{n}(t)$. We also show that the orthogonal polynomials satisfy a second-order ordinary differential equation. By using Dyson’s Coulomb fluid approach, we obtain the large $n$ asymptotic expansions of the recurrence coefficients $\alpha_{n}(t)$ and $\beta_{n}(t)$, the sub-leading coefficient $\mathrm{p}(n,t)$ of the monic orthogonal polynomials, the Hankel determinant $D_{n}(t)$ and the normalized constant $h_{n}(t)$ for fixed $t\in\mathbb{R}^{+}$. We also discuss the long-time asymptotics of these quantities as $t\rightarrow\infty$ for fixed $n\in\mathbb{N}$.
@@ 1 Introduction
with parameters $\alpha>0,\;t>0$ and $\lambda\in\mathbb{R}$.
@@
It is a well-known fact that the Hankel determinant $D_{n}(t)$ can be expressed as the product of the normalized constants for the monic orthogonal polynomials with respect to the weight function (1.1) [12, (2.1.6)]
@@
where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma\leq 2$ and deg $\tau$=1. For semi-classical orthogonal polynomials, they have a weight $w(x)$ that satisfies the Pearson equation (1.4), where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma>2$ or deg $\tau\neq 1$. See, e.g., [23, Section 1.1.1]. It is not difficult to find that (1.1) is a semi-classical weight, since it satisfies the Pearson equation (1.4) with
@@ 2 Ladder operators, compatibility conditions and difference and differential equations
which is the differential equation satisfied by the classical Laguerre polynomials [22, p. 100].
@@
Let $y(t):=1-\frac{1}{1-R_{n}(t)}$, where $R_{n}(t)$ is closely related to the diagonal recurrence coefficient $\alpha_{n}$ by (2.19). Then $y(t)$ satisfies a particular Painlevé V equation [11, p. 1]
$y^{\prime\prime}=\frac{3y-1}{2y(y-1)}(y^{\prime})^{2}-\frac{y^{\prime}}{t}+\frac{(y-1)^{2}}{t^{2}}\left(\mu_{0}y+\frac{\mu_{1}}{y}\right)+\frac{\mu_{2}y}{t}+\frac{\mu_{3}y(y+1)}{y-1},$
@@
Let $\sigma(t):=H_{n}(t)-n\lambda$. Then $\sigma(t)$ satisfies the Jimbo-Miwa-Okamoto $\sigma$-form of the Painlevé V equation [13, (C. 45)]
$(t\sigma^{\prime\prime})^{2}=[\sigma-t\sigma^{\prime}+2(\sigma^{\prime})^{2}+(\nu_{0}+\nu_{1}+\nu_{2}+\nu_{3})\sigma^{\prime}]^{2}-4(\nu_{0}+\sigma^{\prime})(\nu_{1}+\sigma^{\prime})(\nu_{2}+\sigma^{\prime})(\nu_{3}+\sigma^{\prime}),$
@@ 3 Large $n$ asymptotics
Note that there is only a constant difference between $D_{n}(t)$ and $Z_{n}(t)$ using Andréief’s or Heine’s integration formula [22, (2.2.11)]:
@@
According to Dyson’s Coulomb fluid approach [9], the eigenvalues of Hermitian matrices from the unitary ensemble can be approximated as a continuous fluid with the equilibrium density $\sigma(x)$ when $n$ is sufficiently large. When $\alpha\geq 0$, the potential of our problem in (2.7) is either convex or satisfies the condition that $x\mathrm{v}^{\prime}(x)$ increases on $\mathbb{R}^{+}$. In this case, the equilibrium density $\sigma(x)$ is supported on a single interval, denoted by $(0,b)(b>0)$; see Chen and Ismail [3] and also [21, p. 198–199]. According to [3], the equilibrium density $\sigma(x)$ can be determined by the constrained minimization problem:
@@
$+\dfrac{\lambda^{2}t\left[t+2\left({\alpha+\lambda}\right)\right]}{32{n^{3}}}++O(n^{-7/2})$ (3.10)
@@
Multiplying by $\frac{1}{\sqrt{(b-x)x}}$ on both sides of (3.3) and integrating with respect to $x$ from 0 to $b$, and following the similar computations in [17, p. 406–407], we find the Lagrange multiplier $A$ is
@@
$\alpha_{n}=a_{-2}n+a_{-1}n^{1/2}+{a_{0}}+\sum\limits_{j=1}^{\infty}{\frac{a_{j}}{n^{j/2}}},$ (3.16)
${\beta_{n}}=b_{-4}n^{2}+b_{-3}n^{3/2}+b_{-2}n+b_{-1}n^{1/2}+{b_{0}}+\sum\limits_{j=1}^{\infty}{\frac{b_{j}}{n^{j/2}}},$ (3.17)
@@
$\ln D_{n}(t)=$ $\>{n^{2}}\ln n+\left(\alpha+\lambda\right)n\ln n+\dfrac{6\alpha^{2}+6\alpha\lambda+3\lambda^{2}-2}{12}\ln n-\dfrac{3n^{2}}{2}-\tilde{c}_{2}(\alpha,\lambda)n$
$+2\lambda\sqrt{t}\>n^{1/2}-\frac{\lambda t}{2}-\dfrac{\lambda(2\alpha+\lambda)}{4}\ln t-\tilde{c}_{0}(\alpha,\lambda)+\dfrac{\lambda\left[4t^{2}+24t\left(\alpha+\lambda\right)-12\alpha^{2}+3\right]}{48\sqrt{t}\>n^{1/2}}$
@@
where $\tilde{c}_{2}(\alpha,\lambda)$ and $\tilde{c}_{0}(\alpha,\lambda)$ are constants depending only on $\alpha$ and $\lambda$, and
$\tilde{c}_{2}(\alpha,0)=\alpha-\ln(2\pi),$ (3.21)
$\tilde{c}_{0}(\alpha,0)=-\dfrac{\alpha}{2}\ln(2\pi)-2\zeta^{\prime}(-1)+\ln G(\alpha+1),$ (3.22)
@@
be the “free energy”. Chen and Ismail [3] showed that $F_{n}(t)$ can be approximated by the free energy $F[\sigma]$ in (3.2) for sufficiently large $n$. In view of (3), we have the expansion form of $F_{n}(t)$:
${F_{n}}(t)={c_{7}}(t,\alpha,\lambda){n^{2}}\ln n+{c_{6}}(t,\alpha,\lambda)n\ln n+{c_{5}}(t,\alpha,\lambda)\ln n+\sum_{j=-\infty}^{4}{c_{j}(t,\alpha,\lambda)}{n^{j/2}},$ (3.23)
@@
${F_{n}}(t)=$ $-{n^{2}}\ln n-\left(\alpha+\lambda\right)n\ln n-\dfrac{6\alpha^{2}+6\alpha\lambda+3\lambda^{2}-2}{12}\ln n+\dfrac{3n^{2}}{2}+{c_{2}}(t,\alpha,\lambda)n$
$-2\lambda\sqrt{t}\>n^{1/2}+{c_{0}}(t,\alpha,\lambda)-\dfrac{\lambda\left[4t^{2}+24t\left(\alpha+\lambda\right)-12\alpha^{2}+3\right]}{48\sqrt{t}\>n^{1/2}}$
@@
where ${c_{2}}(t,\alpha,\lambda)$ and ${c_{0}}(t,\alpha,\lambda)$ are undetermined. It follows that
$\ln D_{n}(t)=$ $\>{n^{2}}\ln n+\left(\alpha+\lambda\right)n\ln n+\dfrac{6\alpha^{2}+6\alpha\lambda+3\lambda^{2}-2}{12}\ln n-\dfrac{3n^{2}}{2}-{c_{2}}(t,\alpha,\lambda)n$
$+2\lambda\sqrt{t}\>n^{1/2}-{c_{0}}(t,\alpha,\lambda)+\dfrac{\lambda\left[4t^{2}+24t\left(\alpha+\lambda\right)-12\alpha^{2}+3\right]}{48\sqrt{t}\>n^{1/2}}$
$-\dfrac{12\lambda^{2}t^{2}-8t(\alpha+\lambda)\left(4\alpha^{2}+2\alpha\lambda+\lambda^{2}-2\right)+3\lambda^{2}\left(4\alpha^{2}-1\right)}{192tn}+O(n^{-3/2}).$ (3.25)
By taking a derivative of (3) with respect to $t$ and substituting it into (2.31), we find
${H_{n}}(t)=$ $-t\dfrac{d}{dt}{c_{2}}(t,\alpha,\lambda)n+\lambda\sqrt{t}\>n^{1/2}-t\dfrac{d}{dt}{c_{0}}(t,\alpha,\lambda)+\dfrac{\lambda\left[4t^{2}+8t(\alpha+\lambda)+4\alpha^{2}-1\right]}{32\sqrt{t}\>n^{1/2}}$
Comparing (3) with (3.3), we have
$t\frac{d}{dt}{c_{2}}(t,\alpha,\lambda)=0,$
$t\frac{d}{dt}{c_{0}}(t,\alpha,\lambda)=-\dfrac{\lambda\left(2t+2\alpha+\lambda\right)}{4}.$
It follows that
${c_{2}}(t,\alpha,\lambda)=\tilde{c}_{2}(\alpha,\lambda),$ (3.27)
${c_{0}}(t,\alpha,\lambda)=\frac{\lambda t}{2}+\dfrac{\lambda(2\alpha+\lambda)}{4}\ln t+\tilde{c}_{0}(\alpha,\lambda),$ (3.28)
where $\tilde{c}_{2}(\alpha,\lambda)$ and $\tilde{c}_{0}(\alpha,\lambda)$ are constants depending only on $\alpha$ and $\lambda$. Substituting (3.27) and (3.28) into (3) gives the result in (3.4).
If $\lambda=0$, our weight (2.6) is reduced to the classical Laguerre weight $w(x)={x^{\alpha}}{\mathrm{e}^{-x}},\;x\in\mathbb{R}^{+},\alpha>-1$ and the corresponding Hankel determinant has the closed-form expression [16, p. 321]
$D_{n}$ $=\det\left(\int_{0}^{\infty}x^{i+j}{x^{\alpha}}{\mathrm{e}^{-x}}dx\right)_{i,j=0}^{n-1}$
@@
$\ln D_{n}=$ $\>{n^{2}}\ln n+\alpha n\ln n+\dfrac{3\alpha^{2}-1}{6}\ln n-\dfrac{3}{2}{n^{2}}-[\alpha-\ln(2\pi)]n+\dfrac{\alpha}{2}\ln(2\pi)$
$+2\zeta^{\prime}(-1)-\ln G(\alpha+1)+O(n^{-1}).$ (3.29)
The combination of (3.4) and (3) gives the results in (3.21) and (3.22). This completes the proof. ∎
@@
$\ln h_{n}(t)=$  $\>2n\ln n-2n+(1+\alpha+\lambda)\ln n+\alpha+\lambda-\tilde{c}_{2}(\alpha,\lambda)+\dfrac{{\lambda\sqrt{t}}}{n^{1/2}}$
$+\frac{6\alpha(\alpha+\lambda+1)+3\lambda(\lambda+2)+2}{{12n}}+O(n^{-3/2}),$
where $\tilde{c}_{2}(\alpha,\lambda)$ is a constant depending only on $\alpha$ and $\lambda$, and $\tilde{c}_{2}(\alpha,0)$ is given by (3.21).
@@
Remark 5.
It can be seen from Theorem 3.1 that the asymptotic expansions of the recurrence coefficients are singular at the point $t=0$, which leads to the fact that all the other quantities discussed in this section have the same feature. We can not determine the constants $\tilde{c}_{2}(\alpha,\lambda)$ and $\tilde{c}_{0}(\alpha,\lambda)$ completely. This is called the “constant problem” in random matrix theory, which is often a challenging task; cf. [14, 15].

### v2

@@ Abstract
We consider the Hankel determinant and orthogonal polynomials with respect to the deformed Laguerre weight $w(x;t)={x^{\alpha}}{\mathrm{e}^{-x}}{(x+t)^{\lambda}},\;x\in\mathbb{R}^{+}$ with parameters $\alpha>-1,\;t>0$ and $\lambda\in\mathbb{R}$. This problem originates from the information theory of single-user multiple-input multiple-output (MIMO) systems studied by Chen and McKay [IEEE Trans. Inf. Theory 58 (2012) 4594–4634]. By using the ladder operators for orthogonal polynomials with general Laguerre-type weights, we obtain a system of difference equations and a system of differential-difference equations for the recurrence coefficients $\alpha_{n}(t)$ and $\beta_{n}(t)$. We also show that the orthogonal polynomials satisfy a second-order ordinary differential equation. By using Dyson’s Coulomb fluid approach, we obtain the large $n$ asymptotic expansions of the recurrence coefficients $\alpha_{n}(t)$ and $\beta_{n}(t)$, the sub-leading coefficient $\mathrm{p}(n,t)$ of the monic orthogonal polynomials, the Hankel determinant $D_{n}(t)$ and the normalized constant $h_{n}(t)$ for fixed $t\in\mathbb{R}^{+}$. We also discuss the long-time asymptotics of these quantities as $t\rightarrow\infty$ for fixed $n\in\mathbb{N}$. The large $n$ and large $t$ asymptotics of the above quantities are very important for the study of the asymptotics of the mutual information distribution and two fundamental quantities (the outage capacity and the error probability) for single-user MIMO systems.
@@ 1 Introduction
with parameters $\alpha>0,\;t>0$ and $\lambda\in\mathbb{R}$. When $\lambda=0$, the above weight (1.1) is reduced to the classical Laguerre weight $w_{0}(x)={x^{\alpha}}{\mathrm{e}^{-x}},\;x\in\mathbb{R}^{+},$ and the corresponding Hankel determinant is denoted by $D_{n}[w_{0}]$.
@@
It is a well-known fact that the Hankel determinant $D_{n}(t)$ can be expressed as the product of the normalized constants for the monic orthogonal polynomials with respect to the weight function (1.1) [16, (2.1.6)]
@@
where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma\leq 2$ and deg $\tau$=1. For semi-classical orthogonal polynomials, they have a weight $w(x)$ that satisfies the Pearson equation (1.4), where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma>2$ or deg $\tau\neq 1$. See, e.g., [25, Section 1.1.1]. It is not difficult to find that (1.1) is a semi-classical weight, since it satisfies the Pearson equation (1.4) with
@@ 2 Ladder operators, compatibility conditions and difference and differential equations
which is the differential equation satisfied by the classical Laguerre polynomials [24, p. 100].
@@
Let $W_{n}(t):=1-\frac{1}{1-R_{n}(t)}$, where $R_{n}(t)$ is closely related to the diagonal recurrence coefficient $\alpha_{n}$ by (2.19). Then $W_{n}(t)$ satisfies a particular Painlevé V equation [15, p. 1]
$W_{n}^{\prime\prime}=\frac{3W_{n}-1}{2W_{n}(W_{n}-1)}(W_{n}^{\prime})^{2}-\frac{W_{n}^{\prime}}{t}+\frac{(W_{n}-1)^{2}}{t^{2}}\left(\mu_{0}W_{n}+\frac{\mu_{1}}{W_{n}}\right)+\frac{\mu_{2}W_{n}}{t}+\frac{\mu_{3}W_{n}(W_{n}+1)}{W_{n}-1},$
@@
Let $\sigma_{n}(t):=H_{n}(t)-n\lambda$. Then $\sigma_{n}(t)$ satisfies the Jimbo-Miwa-Okamoto $\sigma$-form of the Painlevé V equation [17, (C. 45)]
$(t\sigma_{n}^{\prime\prime})^{2}=[\sigma_{n}-t\sigma_{n}^{\prime}+2(\sigma_{n}^{\prime})^{2}+(\nu_{0}+\nu_{1}+\nu_{2}+\nu_{3})\sigma_{n}^{\prime}]^{2}-4(\nu_{0}+\sigma_{n}^{\prime})(\nu_{1}+\sigma_{n}^{\prime})(\nu_{2}+\sigma_{n}^{\prime})(\nu_{3}+\sigma_{n}^{\prime}),$
@@ 3 Large $n$ asymptotics
Note that there is only a constant difference between $D_{n}(t)$ and $Z_{n}(t)$ using Andréief’s or Heine’s integration formula [24, (2.2.11)]:
@@
According to Dyson’s Coulomb fluid approach [13], the eigenvalues of Hermitian matrices from the unitary ensemble can be approximated as a continuous fluid with the equilibrium density $\sigma(x)$ when $n$ is sufficiently large. When $\alpha\geq 0$, the potential of our problem in (2.7) is either convex or satisfies the condition that $x\mathrm{v}^{\prime}(x)$ increases on $\mathbb{R}^{+}$. In this case, the equilibrium density $\sigma(x)$ is supported on a single interval, denoted by $(0,b)(b>0)$; see Chen and Ismail [5] and also [23, p. 198–199]. According to [5], the equilibrium density $\sigma(x)$ can be determined by the constrained minimization problem:
@@
$+\dfrac{\lambda^{2}t\left[t+2\left({\alpha+\lambda}\right)\right]}{32{n^{3}}}+O(n^{-7/2})$ (3.10)
@@
Multiplying by $\frac{1}{\sqrt{(b-x)x}}$ on both sides of (3.3) and integrating with respect to $x$ from 0 to $b$, and following the similar computations in [19, p. 406–407], we find the Lagrange multiplier $A$ is
@@
$\alpha_{n}\sim a_{-2}n+a_{-1}n^{1/2}+{a_{0}}+\sum\limits_{j=1}^{\infty}{\frac{a_{j}}{n^{j/2}}},$ (3.16)
${\beta_{n}}\sim b_{-4}n^{2}+b_{-3}n^{3/2}+b_{-2}n+b_{-1}n^{1/2}+{b_{0}}+\sum\limits_{j=1}^{\infty}{\frac{b_{j}}{n^{j/2}}},$ (3.17)
@@
$\ln D_{n}(t)=$ $\>{n^{2}}\ln n-\dfrac{3n^{2}}{2}+\left(\alpha+\lambda\right)n\ln n-[\alpha+\lambda-\ln(2\pi)]n+2\lambda\sqrt{t}\>n^{1/2}$
$+\dfrac{6\alpha^{2}+6\alpha\lambda+3\lambda^{2}-2}{12}\ln n-c_{0}+\dfrac{\lambda\left[4t^{2}+24t\left(\alpha+\lambda\right)-12\alpha^{2}+3\right]}{48\sqrt{t}\>n^{1/2}}$
@@
where the constant term reads
$c_{0}=2\zeta^{\prime}(-1)-\ln G(\alpha+1)+\dfrac{\alpha}{2}\ln(2\pi)-\frac{\lambda}{4}\left[2t+\lambda\ln 4+(2\alpha+\lambda)\ln t\right],$
@@
be the “free energy”. Chen and Ismail [5] showed that $F_{n}(t)$ can be approximated by the free energy $F[\sigma]$ in (3.2) for sufficiently large $n$. Chen, Ismail and Van Assche [7] also stated that the approximation is very accurate and effective, and applied the approach to three different examples. More recently, Basor, Chen and Haq [3, Section 7.2] used this approach to derive the large $n$ asymptotic expansion of the Hankel determinant for a modified Jacobi weight and also used other methods to show the results are consistent. In view of (3), we assume that $F_{n}(t)$ has the large $n$ expansion form:
${F_{n}}(t)={c_{7}}{n^{2}}\ln n+{c_{6}}n\ln n+{c_{5}}\ln n+\sum_{j=-\infty}^{4}{c_{j}}{n^{j/2}},$ (3.21)
@@
${F_{n}}(t)=$ $-{n^{2}}\ln n-\left(\alpha+\lambda\right)n\ln n-\dfrac{6\alpha^{2}+6\alpha\lambda+3\lambda^{2}-2}{12}\ln n+\dfrac{3n^{2}}{2}+{c_{2}}n$
$-2\lambda\sqrt{t}\>n^{1/2}+{c_{0}}-\dfrac{\lambda\left[4t^{2}+24t\left(\alpha+\lambda\right)-12\alpha^{2}+3\right]}{48\sqrt{t}\>n^{1/2}}$
@@
where $c_{2}$} and $c_{0}$} are undetermined. It follows that
$\ln D_{n}(t)=$ $\>{n^{2}}\ln n-\dfrac{3n^{2}}{2}+\left(\alpha+\lambda\right)n\ln n-{c_{2}}n+2\lambda\sqrt{t}\>n^{1/2}$
$+\dfrac{6\alpha^{2}+6\alpha\lambda+3\lambda^{2}-2}{12}\ln n-{c_{0}}+\dfrac{\lambda\left[4t^{2}+24t\left(\alpha+\lambda\right)-12\alpha^{2}+3\right]}{48\sqrt{t}\>n^{1/2}}$
$-\dfrac{12\lambda^{2}t^{2}-8t(\alpha+\lambda)\left(4\alpha^{2}+2\alpha\lambda+\lambda^{2}-2\right)+3\lambda^{2}\left(4\alpha^{2}-1\right)}{192tn}+O(n^{-3/2}).$
To determine the two constants $c_{2}$ and $c_{0}$, we use a result from the work of Chen and Lawrence [8]. Following [8], we have the large $n$ leading term asymptotics
$\frac{D_{n}[w]}{D_{n}[w_{0}]}\sim\exp\left(-\lambda\int_{\tilde{a}}^{\tilde{b}}f(x)\sigma_{0}(x)dx-\frac{\lambda}{2}\int_{\tilde{a}}^{\tilde{b}}f(x)\varrho(x)dx\right),$
where
$\sigma_{0}(x)=\frac{\sqrt{(\tilde{b}-x)(x-\tilde{a})}}{2\pi^{2}}P\int_{\tilde{a}}^{\tilde{b}}\frac{\mathrm{v}_{0}^{\prime}(y)dy}{(y-x)\sqrt{(\tilde{b}-y)(y-\tilde{a})}}$
and
$\varrho(x)=\frac{\lambda}{2\pi^{2}\sqrt{(\tilde{b}-x)(x-\tilde{a})}}P\int_{\tilde{a}}^{\tilde{b}}\frac{\sqrt{(\tilde{b}-y)(y-\tilde{a})}}{y-x}f^{\prime}(y)dy,$
and $\mathrm{v}_{0}(x)=-\ln w_{0}(x)=x-\alpha\ln x,\;f(x)=-\ln(x+t),\;x\in\mathbb{R}^{+}$ for our problem. The endpoints $\tilde{a}$ and $\tilde{b}$ are determined by the following two equations:
$\int_{\tilde{a}}^{\tilde{b}}\frac{\mathrm{v}_{0}^{\prime}(x)}{\sqrt{(\tilde{b}-x)(x-\tilde{a})}}dx=0,$
$\frac{1}{2\pi}\int_{\tilde{a}}^{\tilde{b}}\frac{x\>\mathrm{v}_{0}^{\prime}(x)}{\sqrt{(\tilde{b}-x)(x-\tilde{a})}}dx=n.$
It can be found that
$\sigma_{0}(x)=\frac{\sqrt{(\tilde{b}-x)(x-\tilde{a})}}{2\pi x},$
$\varrho(x)=\frac{\lambda}{2\pi\sqrt{(\tilde{b}-x)(x-\tilde{a})}}\left(1-\frac{\sqrt{(\tilde{a}+t)(\tilde{b}+t)}}{x+t}\right),\\$
and $\tilde{a}=2n+\alpha-2\sqrt{n(n+\alpha)},\;\;\tilde{b}=2n+\alpha+2\sqrt{n(n+\alpha)}$. After some elementary computations, we obtain
$\frac{D_{n}[w]}{D_{n}[w_{0}]}\sim$ $\exp\Big[\lambda n\ln n-\lambda n+2\lambda\sqrt{t}\>n^{\frac{1}{2}}+\frac{\lambda}{4}(2\alpha+\lambda)\ln n$
$-\frac{\lambda}{4}\big(2t+\lambda\ln 4+(2\alpha+\lambda)\ln t\big)+o(1)\Big].$ (3.23)
For the classical Laguerre weight $w_{0}(x)={x^{\alpha}}{\mathrm{e}^{-x}},\;x\in\mathbb{R}^{+},\alpha>-1$, it is known that the corresponding Hankel determinant has the closed-form expression [18, p. 321]
$D_{n}[w_{0}]$ $=\det\left(\int_{0}^{\infty}x^{i+j}{x^{\alpha}}{\mathrm{e}^{-x}}dx\right)_{i,j=0}^{n-1}$
@@
$\ln D_{n}[w_{0}]=$ $\>{n^{2}}\ln n-\dfrac{3}{2}{n^{2}}+\alpha n\ln n-[\alpha-\ln(2\pi)]n+\dfrac{3\alpha^{2}-1}{6}\ln n$
$+\dfrac{\alpha}{2}\ln(2\pi)+2\zeta^{\prime}(-1)-\ln G(\alpha+1)+O(n^{-1}).$ (3.24)
The combination of (3) and (3) shows that
$c_{2}=-\alpha-\lambda+\ln(2\pi)$
and
$c_{0}=2\zeta^{\prime}(-1)-\ln G(\alpha+1)+\dfrac{\alpha}{2}\ln(2\pi)-\frac{\lambda}{4}\left[2t+\lambda\ln 4+(2\alpha+\lambda)\ln t\right].$
This completes the proof. ∎
Remark 5.
Recently, Charlier and Gharakhloo [4] studied the large $n$ asymptotics of Hankel determinants for Laguerre-type and Jacobi-type weights. Note that the weight of Charlier and Gharakhloo [4, (1.2)] is a varying weight with the potential $nV(x)$, and our weight in (2.6) is independent of $n$. Therefore, our weight is not a special case of [4, (1.2)] and can not be transformed to their case either. As a result, the large $n$ expansion form of our Hankel determinant in (3.4) is different from [4, Theorem 1.2] since we have the fractional powers of $n$. In addition, it can be seen from Theorem 3.1 that the asymptotic expansions of the recurrence coefficients are singular at the point $t=0$, which leads to the fact that all the other quantities discussed in this section have the same feature.
@@
$\ln h_{n}(t)=$  $\>2n\ln n-2n+(1+\alpha+\lambda)\ln n+\ln(2\pi)+\dfrac{{\lambda\sqrt{t}}}{n^{1/2}}$
$+\frac{6\alpha(\alpha+\lambda+1)+3\lambda(\lambda+2)+2}{{12n}}+O(n^{-3/2}).$
@@
At the end of this section, we would like to mention that the large $n$ asymptotic expansion of the Hankel determinant in Theorem 3.4 can be used to study the asymptotics of the moment generating function of the mutual information for single-user MIMO systems when the number of antennas becomes large. Then, the asymptotics of the outage probability and the upper bound on the error probability can be calculated via the asymptotics of the moment generating function; see [9, Section III].